# The Host Galaxies of Fast Radio Bursts Track a Combination of Stellar Mass and Star Formation, Similar to Type Ia Supernovae

### v1

@@ 3.3 FRB Host-galaxy Sample
We choose to work with the FRB host-galaxy sample presented in K. Sharma et al. (2024), the largest published data set (at the time writing our work) with galaxy properties inferred using Prospector SED-modeling code. As previously discussed, because of systematic differences between galaxy SED-modeling codes, it critical that both the FRB host-galaxy sample and the comparison galaxy sample (in our case the distribution derived by J. Leja et al. 2022) be modeled uniformly using the same code and model assumptions. The FRB host-galaxy sample used here includes 23 hosts localized by ASKAP (A. C. Gordon et al., 2023), 4 hosts found by CHIME (M. Bhardwaj et al., 2024), and 26 hosts from DSA-110 with secure associations (K. Sharma et al., 2024). All of these galaxy’s properties were derived using the Prospector code and follow selection criterion presented in K. Sharma et al. (2024). Following K. Sharma et al. (2024) we also take into account an optical selection bias for galaxies, and generate mock data (see 3.4) that only includes galaxies whose $r$-band magnitude is $\leq 23.5$. We follow the methodology presented in K. Sharma et al. (2024) in estimating the $r$-band magnitude of mock galaxies using a prescribed mass-to-light ratio (K. Sharma et al. 2024; see “Background Galaxy Population” section in that work).
@@ 3.4 Model Testing
Having defined the likelihood $\mathcal{L}$ of a sample of $N$ FRB host galaxies in Equation (3), we are now in a position to test the hypothesis that these $N$ galaxies are drawn from some underlying model distribution. To do so, we use the Monte-Carlo sampling code emcee (D. Foreman-Mackey et al., 2013) to randomly draw sets of $N$ galaxies from the weighted distribution function describing a given model. Specifically, we generate sets of $N=51$ galaxies such that the $i^{\rm th}$ galaxy in this set (where $i=1,...,N$) is drawn from the weighted probability distribution function $\propto W(M,{\rm SFR})\rho\left(M,{\rm SFR},|z_{i}\right)$, where $z_{i}$ is the redshift of the $i^{\rm th}$ galaxy in the comparison FRB host-galaxy sample. We repeat this process 32,000 times, effectively generating a large number of mock data-sets (each with $N=51$ galaxies) that can be compared to the observed data. The choice of 32,000 mock data sets is arbitrary, but chosen to ensure sufficiently large statistics.
@@
where $\mathcal{L}_{\rm data}$ is the likelihood of the observed FRB host-galaxy sample within this model. The p-value defined via Equation (8) can be interpreted as the probability that a set of $N$ galaxies randomly drawn from the assumed model distribution would have a likelihood as low, or lower than that of the observed galaxy sample. For example, a p-value of $p=0.05$ implies that there is only a 5% chance that galaxies drawn from the given model distribution would have a likelihood as extreme as the FRB host-galaxy sample. In following standard conventions, we adopt $p<0.05$ as a threshold p-value below which we consider a given model to be ruled out by the data (in the sense that there would only be a $<5\%$ probability of observing the sample of known FRB host-galaxies if the model were correct).
@@ 4 Results and implications
Table 1: The p-values of different models for FRB host-galaxies, as obtained by our methodology. The ‘conservative’ sub-sample considers only FRB host-galaxies that do not require any extension of the distribution (see §3.3).
@@
footnotetext: aMass-weighted model (Equation 4); bSFR-weighted model (Equation 5); cGlobular-cluster-weighted model (Equation 7); dSFR+Z model defined via Equation (9) and following K. Sharma et al. (2024); eThe 90% upper confidence limit on the p-value considering the quoted uncertainties in inferred FRB host-galaxy masses and SFRs; fThe nominal p-value obtained for the quoted (median) masses and SFRs of each FRB host galaxy. gApplied to the subsample of FRB host-galaxies with $z<0.2$ ($N=20$).
Table 1 summarizes the resulting p-values for the mass-weighted, SFR-weighted, and GC-weighted models. Using our full sample of $N=51$ host galaxies, we can rule out the SFR and GC models at a significance of $p\sim 0.02$, while the mass-weighted model cannot be ruled out ($p\sim 0.07$). This implies that FRBs do not linearly track SFR in the Universe but are potentially consistent with following stellar mass, though only marginally so (as we discuss in §4.2, we disfavor this scenario). The results also show that FRBs, as a whole population, are inconsistent with tracking the mass of GCs. Considering a conservative sub-sample of $N=23$ FRBs whose host galaxies do not require any extension of the J. Leja et al. (2022) distribution function (see §3.3 and Tables 2,3), we cannot rule out any of the three models, although the p-values of both SFR and GC models are only marginally consistent with the data (given our predetermined threshold of $p=0.05$).
To the best of our knowledge this is the first time that it has been explicitly shown that FRBs are inconsistent with a GC-origin, as a whole population. This may not be a very surprising conclusion, but it is important given the identification of at least one FRB to a GC (F. Kirsten et al., 2022). The relative fraction of FRB sources that could be located in GCs remains an open question. Our analysis uses host-galaxy information to show that it is statistically unlikely that all FRBs originate from GCs. But a more realistic model whereby only a fraction of FRBs are associated with GCs, would likely be consistent with the current host-galaxy data.
@@
[…]
where $M_{Z}$ is a cutoff mass (denoted $M_{c}$ in K. Sharma et al. 2024) that accounts for the fact that galaxies with lower mass (and hence, on average, lower metallicity) do not produce FRBs. The parameter $\beta$ controls the steepness of this cutoff. We use the values obtained by K. Sharma et al. (2024) for a subsample of FRB host galaxies at low redshifts, $z<0.2$: $\log_{10}(M_{Z}/M_{\odot})=8.98$, and $\beta\simeq 37,276$. For such large $\beta$, the sigmoid in the weighting-function (the second term in Equation 9) essentially becomes a step-function, selecting only galaxies with masses greater than $M_{Z}$.
@@ 4.1 Mixed-Population Model
We continue by investigating a mixed model in which FRBs track a linear combination of both stellar mass and SFR, as described via Equation (6). Figure 3 shows the resulting p-values for such a model, as a function of $A/B$—the ratio of the stellar mass prefactor and the SFR prefactor in Equation (6). Orange (blue) points show $p_{\rm nom}$ ($p_{90}$) for different values of $A/B$. The right-hand side of this plot corresponds to the purely mass-weighted model, while the left-hand side corresponds to the SFR-weighted scenario. The p-values in either case are consistent with the values quoted in Table 1 for those two models. The SFR-weighted end falls below our threshold $p=0.05$ value (horizontal dashed-grey curve) implying that this models is ruled out, as previously discussed. The mass-weighted end asymptotes to a value slightly above $p=0.05$ and can cannot be strictly ruled out by this analysis, however there is only marginal support for such a model (see also §4.2). Alternatively, Figure 3 shows that a mixed model with $A/B\sim 10^{-10}\,{\rm yr}^{-1}$ seems to be preferred by the current data. This suggests that FRBs could track a combination of both stellar mass and star formation.
@@
The preferred value of $A/B\approx 3\times 10^{-11}\,{\rm yr}^{-1}$ that we find here is surprisingly close to the value inferred for Type Ia SNe, $A/B\approx 1.7^{+2.3}_{-0.9}\times 10^{-11}\,{\rm yr}^{-1}$ (E. Scannapieco & L. Bildsten, 2005), although note that there are significant uncertainties associated with the exact value of our preferred $A/B$. This implies that the progenitors of FRBs and Type Ia SNe may be related. In particular, this aligns with the idea that FRBs might be associated with accretion-induced collapse of a white-dwarf, as such events are thought to form through similar channels as Type Ia SNe (K. Kashiyama & K. Murase, 2017; E. Waxman, 2017; B. Margalit et al., 2019; T. Eftekhari et al., 2025).
@@
One interpretation of the mixed model is that there are two distinct formation channels (components) for FRB sources: one that follows star-formation and one that is dependent on stellar mass. Viewed this way, the parameter $A/B$ can be interpreted as a specific star-formation rate ${\rm sSFR}\equiv{\rm SFR}/M$ that delineates between the two components. FRB sources located in host-galaxies whose specific SFR exceeds ${\rm sSFR}>A/B\approx 10^{-10.5}\,{\rm yr}^{-1}\approx 0.03\,{\rm Gyr}^{-1}$ are likely to have formed through the star-formation channel, while sources in galaxies with ${\rm sSFR}<A/B$ are more likely to have formed via the mass-dependent channel. Another interpretation of the mixed model is that FRB sources are born through a single channel, but one that has an extended time-delay distribution (meaning, a fraction of sources are formed long after their progenitor stars were born). Although the model does not explicitly account for the delay-time distribution, the combination of stellar-mass and SFR can mimic this behavior. This is the common interpretation for short GRBs and Type Ia SNe (e.g., E. Scannapieco & L. Bildsten, 2005; W. Fong et al., 2013). The extended delay time in these settings is then typically associated with the gravitational-inspiral time of a compact object binary.
[…]
$\frac{\left.\mathcal{R}_{\rm FRB}(t)\right|_{M}}{\left.\mathcal{R}_{\rm FRB}(t)\right|_{\rm SFR}}=\frac{A}{B}\tau\left(e^{t/\tau}-1\right)\underset{t=10\,{\rm Gyr}}{\sim}9.$ (10)
@@
That is, after $t=10\,{\rm Gyr}$, only $\sim$10(one out of 10) of FRB sources are born via channels that track SFR. Conversely, the ratio of the total number of FRB sources born through either channel throughout the lifetime of such a galaxy is
$\frac{\left.N_{\rm FRB}(t)\right|_{M}}{\left.N_{\rm FRB}(t)\right|_{\rm SFR}}=\frac{A}{B}\tau\frac{(t/\tau)e^{t/\tau}+1}{e^{t/\tau}-1}\underset{t=10\,{\rm Gyr}}{\sim}0.3.$ (11)
@@
This implies that over the lifetime of a galaxy roughly $\sim$80(one out of 1.3) of FRB sources will be born through SFR-related channels. Of course, it is important to bear in mind the order-of-magnitude uncertainty in the allowed $A/B$ values, as well as the grossly over-simplified galaxy growth model that goes into these estimates. They are nonetheless illustrative and useful as a comparison to Type Ia SNe progenitors.
@@
Using our preferred mixed model we can also predict the fraction of quiescent FRB host galaxies. We simulate a mock population of galaxies drawn from the weighted distribution defined by Equation (6) with $A/B=10^{-10.5}\,{\rm yr}^{-1}$ at redshift $z=0.3$. We then classify galaxies as quiescent, transitioning, or star-forming following the methodology described in A. C. Gordon et al. (2023). Specifically, we use the mass-doubling number $\mathcal{D}={\rm sSFR}\times t_{\rm H}(z)$, where $t_{\rm H}(z)$ is the Hubble time at redshift $z$, and classify galaxies as quiescent if $\mathcal{D}<1/20$, transitioning if $31/20<\mathcal{D}<1/$ 3, and star-forming if $\mathcal{D}>1/3$ (S. Tacchella et al., 2022). According to this criterion we find that 12% of the simulated galaxies are quiescent, 10% are transitioning, and 78% are star-forming at $z=0.3$. At redshift $z=0.5$ the associated breakdown is 9% quiescent, 7% transitioning, and 84% star-forming.
@@
[…]
is a redshift-specific normalization that can be interpreted as the ratio between the cosmic star-formation density $\Psi(z)$ to the cosmic stellar-mass density $\rho_{*}(z)$ at redshift $z$ Within this framework our preferred value of $A/B=10^{-10.5}\,{\rm yr}^{-1}$ translates into $\chi_{\rm SFR}\approx 0.65$ at redshift $z=0.1$, $\chi_{\rm SFR}\approx 0.74$ at $z=0.3$, and $\chi_{\rm SFR}\approx 0.80$ at $z=0.5$. These values are all consistent with the constraints on $\chi_{\rm SFR}$ obtained by N. Loudas et al. (2025).

### v2

@@ 3.3 FRB Host-galaxy Sample
We choose to work with the FRB host-galaxy samples presented in A. C. Gordon et al. (2023), M. Bhardwaj et al. (2024), and K. Sharma et al. (2024), the largest published data set (at the time writing our work) with galaxy properties inferred using Prospector SED-modeling code. As previously discussed, because of systematic differences between galaxy SED-modeling codes, it critical that both the FRB host-galaxy sample and the comparison galaxy sample (in our case the distribution derived by J. Leja et al. 2022) be modeled uniformly using the same code and model assumptions. The FRB host-galaxy sample used here includes 23 hosts localized by ASKAP (A. C. Gordon et al., 2023), 4 hosts found by CHIME (M. Bhardwaj et al., 2024), and 26 hosts from DSA-110 with secure associations (K. Sharma et al., 2024). All of these galaxy’s properties were derived using the Prospector code and follow selection criterion presented in K. Sharma et al. (2024). Following K. Sharma et al. (2024) we also take into account an optical selection bias for galaxies, and generate mock data (see 3.4) that only includes galaxies whose $r$-band magnitude is $\leq 23.5$. We follow the methodology presented in K. Sharma et al. (2024) in estimating the $r$-band magnitude of mock galaxies using a prescribed mass-to-light ratio (K. Sharma et al. 2024; see “Background Galaxy Population” section in that work).
@@ 3.4 Model Testing
Having defined the likelihood $\mathcal{L}$ of a sample of $N$ FRB host galaxies in Equation (3), we are now in a position to test the hypothesis that these $N$ galaxies are drawn from some underlying model distribution. To do so, we use the Monte-Carlo sampling code emcee (D. Foreman-Mackey et al., 2013) to randomly draw sets of $N$ galaxies from the weighted distribution function describing a given model. Specifically, we generate sets of $N=51$ galaxies such that the $i^{\rm th}$ galaxy in this set (where i=1,,Ni=1,...,N) is drawn from the weighted probability distribution function $\propto W(M,{\rm SFR})\rho\left(M,{\rm SFR},|z_{i}\right)$, where $z_{i}$ is the redshift of the $i^{\rm th}$ galaxy in the comparison FRB host-galaxy sample. We repeat this process 32,000 times, effectively generating a large number of mock data-sets (each with $N=51$ galaxies) that can be compared to the observed data. The choice of 32,000 mock data sets is arbitrary, but chosen to ensure sufficiently large statistics.
@@
where $\mathcal{L}_{\rm data}$ is the likelihood of the observed FRB host-galaxy sample within this model. The p-value defined via Equation (8) can be interpreted as the probability that a set of $N$ galaxies randomly drawn from the assumed model distribution would have a likelihood as low, or lower than that of the observed galaxy sample. For example, a p-value of $p=0.05$ implies that there is only a 5% chance that galaxies drawn from the given model distribution would have a likelihood as extreme as the FRB host-galaxy sample. In following standard conventions, we adopt $p<0.05$ as a threshold p-value below which we consider a given model to be ruled out by the data (in the sense that there would only be a $5\%$ probability of observing the sample of known FRB host-galaxies if the model were correct).
@@ 4 Results and implications
Table 1: The p-values of different models for FRB host-galaxies, as obtained by our methodology. The ‘conservative’ sub-sample considers only FRB host-galaxies that do not require any extension of the distribution (see §3.3). footnotetext: aMass-weighted model (Equation 4); bSFR-weighted model (Equation 5); cGlobular-cluster-weighted model (Equation 7); dSFR+Z model defined via Equation (9) and following K. Sharma et al. (2024); eThe 90% upper confidence limit on the p-value considering the quoted uncertainties in inferred FRB host-galaxy masses and SFRs; fThe nominal p-value obtained for the quoted (median) masses and SFRs of each FRB host galaxy. gApplied to the subsample of FRB host-galaxies with $z<0.2$ ($N=20$).
@@
Table 1 summarizes the resulting p-values for the mass-weighted, SFR-weighted, and GC-weighted models. Using our full sample of $N=51$ host galaxies, we can rule out the SFR and GC models at a significance of $p\approx 0.01$, $p\approx 0.03$ respectfully, while the mass-weighted model is only marginally ruled out ($p\approx 0.05$). This implies that FRBs do not linearly track SFR in the Universe but are potentially consistent with following stellar mass, though only marginally so (as we discuss in §4.2, we disfavor this scenario). The results also show that FRBs, as a whole population, are inconsistent with tracking the mass of GCs. Considering a conservative sub-sample of $N=23$ FRBs whose host galaxies do not require any extension of the J. Leja et al. (2022) distribution function (see §3.3 and Tables 2,3), we cannot rule out any of the three models, although the p-values of both SFR and GC models are only marginally consistent with the data (given our predetermined threshold of $p=0.05$).
A. Rao et al. (2025) have recently used GC simulations to show that NSs born through binary white-dwarf coalescence in GCs can contribute at most $\lesssim 1\%$ to the formation rate of FRB sources at $z\lesssim 1$. Our results add to these findings using alternative methods, and support the conclusion that FRBs are inconsistent with a GC-origin as a whole population. This may not be a very surprising conclusion, but it is important given the identification of at least one FRB to a GC (F. Kirsten et al., 2022). The relative fraction of FRB sources that could be located in GCs remains an open question. Our analysis uses host-galaxy information to show that it is statistically unlikely that all FRBs originate from GCs. But a more realistic model whereby only a fraction of FRBs are associated with GCs, would likely be consistent with the current host-galaxy data (e.g., A. Rao et al. 2025 find that $\lesssim 1\%$ of FRBs could arise from GCs).
@@
$\text{`SFR+Z':}~~~~W_{i}={\rm SFR}_{i}\times\left[1+(M_{i}/M_{Z})^{-\beta}\right]^{-1},$ (9)
where $M_{Z}$ is a cutoff mass (denoted $M_{c}$ in K. Sharma et al. 2024) that accounts for the fact that galaxies with lower mass (and hence, on average, lower metallicity) do not produce FRBs. The parameter $\beta$ controls the steepness of this cutoff. We use the values obtained by K. Sharma et al. (2024) for a subsample of FRB host galaxies at low redshifts, $z<0.2$: $\log_{10}(M_{Z}/M_{\odot})=8.98$, and $\beta\simeq 37,276$. For such large $\beta$, the sigmoid in the weighting-function (the second term in Equation 9) essentially becomes a step-function, selecting only galaxies with masses greater than $M_{Z}$. According to the mass-metallicity relations adopted in K. Sharma et al. (2024), this threshold mass corresponds to a threshold oxygen abundance of $12+\log({\rm O}/{\rm H})\sim 8$ and a metallicity of $\log(Z/Z_{\odot})\sim-0.6$.
@@ 4.1 Mixed-Population Model
We continue by investigating a mixed model in which FRBs track a linear combination of both stellar mass and SFR, as described via Equation (6). Figure 3 shows the resulting p-values for such a model, as a function of A/BA/B—the ratio of the stellar mass prefactor and the SFR prefactor in Equation (6). Orange (blue) points show $p_{\rm nom}$ ($p_{90}$) for different values of $A/B$. The right-hand side of this plot corresponds to the purely mass-weighted model, while the left-hand side corresponds to the SFR-weighted scenario. The p-values in either case are consistent with the values quoted in Table 1 for those two models. The SFR-weighted end falls below our threshold $p=0.05$ value (horizontal dashed-grey curve) implying that this models is ruled out, as previously discussed. The mass-weighted end asymptotes to a value slightly above $p=0.05$ and can cannot be strictly ruled out by this analysis, however there is only marginal support for such a model (see also §4.2). Alternatively, Figure 3 shows that a mixed model with $A/B\sim 10^{-10}\,{\rm yr}^{-1}$ seems to be preferred by the current data. This suggests that FRBs could track a combination of both stellar mass and star formation.
@@
The preferred value of $A/B\approx 3\times 10^{-11}\,{\rm yr}^{-1}$ that we find here is close to the value inferred for Type Ia SNe, $A/B\approx 1.7^{+2.3}_{-0.9}\times 10^{-11}\,{\rm yr}^{-1}$ (E. Scannapieco & L. Bildsten, 2005). This implies that the progenitors of FRBs and Type Ia SNe may be related. In particular, this aligns with the idea that FRBs might be associated with accretion-induced collapse of a white-dwarf, as such events are thought to form through similar channels as Type Ia SNe (K. Kashiyama & K. Murase, 2017; E. Waxman, 2017; B. Margalit et al., 2019; T. Eftekhari et al., 2025). We note however that there are significant uncertainties associated with our preferred $A/B$ value, and therefore caution is warranted in interpreting these results. As mentioned in §3, using a different SED modeling code can potentially affect the specific star-formation rate of galaxies ${\rm sSFR}\equiv{\rm SFR}/M$, and hence also the inferred A/B values, by up to an order of magnitude (e.g., J. Leja et al., 2020). Although the A/B values found by E. Scannapieco & L. Bildsten (2005) are not obtained using Prospector, their methodology differs substantially from our present work, and does not rely on detailed galaxy SED fitting. Instead, E. Scannapieco & L. Bildsten (2005) use a sub-sample of host galaxies that are categorized as either “star-forming” or “quiescent”, and determine the A and B values by calibrating to the observed CCSN rate in these galaxies. Furthermore, recent work by A. E. Nugent et al. (2025) re-analyzed Type Ia SNe host galaxies uniformly using Prospector, and found that the A/B values inferred by E. Scannapieco & L. Bildsten (2005) are consistent with those inferred using this SED-fitting code.
@@
Proceeding with these caveats in mind, we consider some implications of the preferred value of $A/B\approx 10^{-10.5}\,{\rm yr}^{-1}$ that we find for FRBs. One interpretation of the mixed model is that there are two distinct formation channels (components) for FRB sources: one that follows star-formation and one that is dependent on stellar mass. Viewed this way, the parameter $A/B$ can be interpreted as a $\rm sSFR$} that delineates between the two components. FRB sources located in host-galaxies whose specific SFR exceeds ${\rm sSFR}>A/B\approx 10^{-10.5}\,{\rm yr}^{-1}\approx 0.03\,{\rm Gyr}^{-1}$ are likely to have formed through the star-formation channel, while sources in galaxies with ${\rm sSFR}<A/B$ are more likely to have formed via the mass-dependent channel. Another interpretation of the mixed model is that FRB sources are born through a single channel, but one that has an extended time-delay distribution (meaning, a fraction of sources are formed long after their progenitor stars were born). Although the model does not explicitly account for the delay-time distribution, the combination of stellar-mass and SFR can mimic this behavior. This is the common interpretation for short GRBs and Type Ia SNe (e.g., E. Scannapieco & L. Bildsten, 2005; W. Fong et al., 2013). The extended delay time in these settings is then typically associated with the gravitational-inspiral time of a compact object binary.
Motivated by the similarity with Type Ia SNe host-galaxies, we follow the analysis of E. Scannapieco & L. Bildsten (2005) and adopt a toy-model for galaxy formation whereby stellar mass grows only as a consequence of star-formation such that $M(t)=\int{\rm SFR}(t)dt$ and taking an exponential star-formation history, ${\rm SFR}(t)={\rm SFR}_{0}e^{-t/\tau}$ with $\tau=2\,{\rm Gyr}$ (this neglects galaxy mergers and hierarchical structure formation).
Using this simplified approach, we investigate implications of our preferred mixed-population model. We assume that FRBs track a combination of both stellar mass and star-formation, with $A/B=10^{-10.5}\,{\rm yr}^{-1}$ (Equation 6). The birth-rate of FRB sources originating from the stellar-mass component at some time $t$, $\left.\mathcal{R}_{\rm FRB}(t)\right|_{A}$, can be compared to the rate of sources tracking the SFR component, $\left.\mathcal{R}_{\rm FRB}(t)\right|_{B}$, where $\mathcal{R}_{\rm FRB}=\left.\mathcal{R}_{\rm FRB}\right|_{A}+\left.\mathcal{R}_{\rm FRB}\right|_{B}$ is the total formation rate of FRB sources. The relative rate in this scenario is
$\frac{\left.\mathcal{R}_{\rm FRB}(t)\right|_{A}}{\left.\mathcal{R}_{\rm FRB}(t)\right|_{B}}=\frac{A}{B}\tau\left(e^{t/\tau}-1\right)\underset{t=10\,{\rm Gyr}}{\sim}9.$ (10)
@@
That is, after $t=10\,{\rm Gyr}$, only $\sim$10% (one out of 10) of FRB sources are born via channels that track SFR. Conversely, the ratio of the total number of FRB sources born through either channel throughout the lifetime of such a galaxy is
$\frac{\left.N_{\rm FRB}(t)\right|_{A}}{\left.N_{\rm FRB}(t)\right|_{B}}=\frac{A}{B}\tau\frac{(t/\tau)e^{t/\tau}+1}{e^{t/\tau}-1}\underset{t=10\,{\rm Gyr}}{\sim}0.3.$ (11)
@@
This implies that over the lifetime of a galaxy roughly $\sim$80% (one out of 1.3) of FRB sources will be born through SFR-related channels. Of course, it is important to bear in mind the order-of-magnitude uncertainty in the allowed $A/B$ values, as well as the grossly over-simplified galaxy growth model that goes into these estimates. They are nonetheless illustrative and useful as a comparison to Type Ia SNe progenitors.
@@
Using our preferred mixed model we can also predict the fraction of quiescent FRB host galaxies. We simulate a mock population of galaxies drawn from the weighted distribution defined by Equation (6) with $A/B=10^{-10.5}\,{\rm yr}^{-1}$ at redshift $z=0.3$. We then classify galaxies as quiescent, transitioning, or star-forming following the methodology described in A. C. Gordon et al. (2023). Specifically, we use the mass-doubling number $\mathcal{D}={\rm sSFR}\times t_{\rm H}(z)$, where $t_{\rm H}(z)$ is the Hubble time at redshift $z$, and classify galaxies as quiescent if $\mathcal{D}<1/20$, transitioning if $1/20<\mathcal{D}<1/3$, and star-forming if $\mathcal{D}>1/3$ (S. Tacchella et al., 2022). According to this criterion we find that 13% of the simulated galaxies are quiescent, 10% are transitioning, and 77% are star-forming at $z=0.3$. At redshift $z=0.5$ the associated breakdown is 10% quiescent, 8% transitioning, and 82% star-forming.
@@
We note again that our preferred A/B value can be interpreted as a sSFR. The sSFR of a “typical” galaxy can be related to the inverse of the Hubble time, since the time-averaged SFR over the history of the galaxy roughly satisfies $\langle{\rm SFR}\rangle_{t}\times t_{\rm H}\sim M$. Our preferred A/B $\sim 10^{-10}\,{\rm yr}^{-1}$ value is therefore consistent with the average sSFR expected for a “typical” low-redshift galaxy, at least at the order-of-magnitude level. This can be interpreted as a consequence of FRB sources being found in a wide range of galaxy types. If all FRBs were instead formed in extreme starburst galaxies (or alternatively, in highly-quiescent galaxies), their sSFRs would be much higher (lower) than $t_{\rm H}^{-1}$ and the inferred A/B values would differ markedly from the inverse Hubble time.
@@
$N(z)=\frac{\int{\rm SFR}\rho\left(M,{\rm SFR}|z\right)\,d{\rm SFR}\,dM}{\int M\rho\left(M,{\rm SFR}|z\right)\,d{\rm SFR}\,dM}$ (13)
is a redshift-specific normalization that can be interpreted as the ratio between the cosmic star-formation density $\Psi(z)$ to the cosmic stellar-mass density $\rho_{*}(z)$ at redshift $z$ . Within this framework our preferred value of $A/B=10^{-10.5}\,{\rm yr}^{-1}$ translates into $\chi_{\rm SFR}\approx 0.65$ at redshift $z=0.1$, $\chi_{\rm SFR}\approx 0.74$ at $z=0.3$, and $\chi_{\rm SFR}\approx 0.80$ at $z=0.5$. These values are all consistent with the constraints on $\chi_{\rm SFR}$ obtained by N. Loudas et al. (2025).